# Foundational theories of hesitant fuzzy sets and hesitant fuzzy information systems and their applications for multi-strength intelligent classifiers

### v3

@@ 6 Conclusion
The definition of inclusion relationship is one of the most foundational concepts of sets. Based on the hesitant fuzzy membership degree of discrete form, we propose several kinds of inclusion relationships for hesitant fuzzy sets. Following the definitions of several kinds of inclusion relationships, some foundational propositions of hesitant fuzzy sets and the families of hesitant fuzzy sets are investigated. Some rules hold in classical sets but not in hesitant fuzzy sets are also shown. In subsequent studies of hesitant fuzzy sets, researchers must avoid the intuitions of classical sets and cannot use these rules.
Furthermore, we investigate some foundational theorems of hesitant fuzzy information systems with respect to parameter reductions and present an example and an algorithm in hesitant fuzzy information systems to illustrate the processes of parameter reductions. Finally, we propose a multi-strength intelligent classifier to make health state diagnoses for complex systems.
In our future works, we will investigate more applications of the multi-strength intelligent classifiers, saturation reductions in hesitant fuzzy information systems, and foundational theories of knowledge bases with different strengths of knowledge, among other topics.
@@ Acknowledgments
The authors thank the editors and the anonymous reviewers for their helpful comments and suggestions that have led to this improved version of the paper. The work was supported by the National Natural Science Foundations of China (No. 71904043).

### v4

@@ 4 Conclusion
The definition of the inclusion relationship is one of the most foundational concepts of sets. Based on the discrete form of hesitant fuzzy membership degrees, we propose several kinds of inclusion relationships for hesitant fuzzy sets. Furthermore, we present some propositions of hesitant fuzzy sets and some propositions of the families of hesitant fuzzy sets based on the proposed inclusion relationships. Some rules that hold in classical sets do not apply to hesitant fuzzy sets, as shown in Section 2. In subsequent studies of hesitant fuzzy sets, researchers must avoid the intuitive understanding of classical sets and cannot use these rules.
In our future work, we will investigate more applications of multi-strength intelligent classifiers, saturation reductions in HFISs, and foundational theories of knowledge bases with varying strengths of knowledge, among other topics.
@@ Acknowledgments
The authors thank the editors and the anonymous reviewers for their helpful comments and suggestions that have led to this improved version of the paper. This work was supported by the National Social Science Foundation Youth Project of China (No. 24CTQ029).